a rectangle has perimeter=26 find diagonal
Answer by Alan3354(69443)

a rectangle has perimeter=26 find diagonal
---------------
That's not sufficient info.
The diagonal is a minumum is it's a square:
d^2 = 6.5^2 + 6.5^2
d^2 = 82.5
d = ~ 9.192
-----------
If it's 10 by 3:
d^2 = 109
d = 10.44
----------
The diagonal is a max when the rectangle is 13- (just under 13, 12.99999...) by 0.000000...1. Then the diagonal is 13-